To neutralize 100 g of HNO3 solution, 30 cm3 of 15% NaOH solution with a density of 1.15 g / cm3 was used. Calculate the mass fraction (in%) of nitric acid in the solution!


1
Expert's answer
2021-05-17T04:20:18-0400

Let the mass of the total solution =100 gm

So mass of the "NaOH" becomes=15 gm {That is 15% of the mass of the solution}


"d_{solution}=\dfrac{m_{solution}}{V_{solution}}"


"V_{solution}=\dfrac{100}{1.15} =86.95ml"


"n_{NaOH}=\dfrac{15}{40}=0.375moles"


Molarity of NaOH "=" "\dfrac{n_{NaOH}}{V_{solution}}\times1000=\dfrac{0.375}{86.95}\times1000=4.31M"


Moles of NaOH required "=4.31\times \dfrac{30}{1000}=0.129 moles"


Moles of "HNO_3" needed "(n_{HNO_3})" "=0.129moles"


"n_{HNO_3}=\dfrac{m_{HNO_3}}{molarmass}"


Mass of "HNO_3=0.129\times63=8.127gm"


Now mass fraction of "HNO_3=\dfrac{8.127}{100}\times100=8.127" %